OpenStudy (keana):

What is the y coordinate of the y-intercept of the line that passes through the points (-4, -1) and (4, 5)?

OpenStudy (ghazi):

(y+1)= (5+1)/(4+4)*(x+4)...put x=0 and find y

OpenStudy (ghazi):

\[(y+1)=\frac{ 6 }{ 8 }(x+4)\]

OpenStudy (keana):

ok we got that but the answer is not what we are able to find, we can't get there

OpenStudy (ghazi):

did you get y=2

OpenStudy (ghazi):

\[y+1=\frac{ 6 }{ 8 }*4=3\]

OpenStudy (ghazi):

y+1=3, y=2

OpenStudy (keana):

nope

OpenStudy (ghazi):

it must be 2